# Event Topology Classifiers at the Large Hadron Collider

### v1

@@ II Event Classifiers
In this section, we define different event shape observables used at the LHC energies. In addition, the correlation of these event shape observables with the number of multi-partonic interactions ($N_{\rm mpi}$) is shown and discussed.
II.1 Different types of events: elastic and inelastic collisions
@@
The final state particle production in hadronic collisions, such as proton-on-proton, is a probe for the initial interactions among the colliding hadrons. An interaction can be classified as elastic or inelastic. In elastic collisions, the incoming and outgoing protons are the same, which takes place via exchange of Pomeron [^1] Pomeron is a colour singlet particle that dominates the elastic scatterings at high energies and has the quantum numbers of the vacuum. [85]. In contrast, in inelastic collisions, the final state particles are not identical to the initial colliding protons. An inelastic event can be diffractive, which is signified by a small energy exchange between the protons. Sometimes, one or both protons interact with the Pomerons and dissociate into multiple final-state particles. The event is single diffractive (SD) if only one proton dissociates while the other one survives as it is ($p+p\rightarrow p+X$). Similarly, when both the protons dissociate to form multiple particles in the final state, it is called a double diffractive (DD) event ($p+p\rightarrow X_{1}+X_{2}$). Sometimes, two Pomerons are exchanged and both the protons survive along with the production of new particles, i.e., $p+p\rightarrow p+p+X$, it is called a central diffractive (CD) event. Additionally, when more than two Pomerons are exchanged, there is a possibility for both protons to dissociate and lead to the production of new particles. These events are usually accompanied by large particle multiplicity in the final state as compared to other diffractive events. Figure 3 shows the pictorial representation of different types of possible interactions in $pp$ collisions. The inelastic events without SD are called Non-Single Diffractive (NSD). At the LHC, the measurements are usually performed for NSD or inelastic (INEL) events.
II.2 Impact parameter and centrality
@@
Figure 4: A schematic representation of the correlation between the final state charged-particle multiplicity distribution ($N_{\rm ch}$) with the Glauber calculated quantities ($b$, and $N_{\rm part}$). The collision centrality classification in experiments is shown explicitly. Figure taken from Ref. [86]
@@
The impact parameter ($b$) of a collision is one of the key parameters that substantially influence the particle production in the final state. However, its measurement is not trivial in experiments as the length scale of the impact parameter ranges in the level of a few fermis. Thus, experimental data are typically categorised by the centrality, defined as percentile of events obtained from the number of produced particles or by the number of participant nucleons, $N_{\rm part}$ that is registered in detectors through the estimation of number of spectator nucleons via $N_{\rm part}=2A-N_{\rm spec}$, where $A$ is the mass number and $N_{\rm spec}$ is the number of spectator nucleons. For example, in ALICE experiment, $N_{\rm spec}$ is given by $E_{\rm ZDC}/E_{\rm A}$, where, $E_{\rm ZDC}$ is the energy deposited in zero-degree calorimeters (ZDC) and $E_{\rm A}$ is beam energy per nucleon. Centrality allows for dividing events into different classes. However, phenomenologically, one needs to assign an impact parameter to a given centrality as depicted in Fig. 4. Thus, theoretical techniques, using the Glauber formalism [86, 87, 88, 89] have been developed to allow the estimation of impact parameter along with the number of participant nucleons from experimental data, which considers multiple scattering of nucleons in nuclear targets. Here, we briefly explain how the total inelastic cross-section, the number of binary collisions, and the number of participants depend on the impact parameter.
@@
Figure 5: Schematic representation of the Optical Glauber Model geometry, with transverse (a) and longitudinal (b) views. Figure taken from Ref. [86]
@@
For a collision of two heavy-nuclei, $A$ and $B$ at relativistic speeds with impact parameter ${\bf b}$ as shown in Fig. 5, the inelastic cross-section can be defined as
$\sigma^{\rm inel}_{AB}({\bf b})$ $=$ $\int 2\pi\rm b\;d{\bf b}\left[1-\left(1-T_{AB}({\bf b})\sigma_{NN}^{\rm inel}\right)^{AB}\right]$ (1)
$\simeq$ $\int 2\pi\rm b\;\;d{\bf b}\left[1-\exp\left[-ABT_{AB}({\bf b})\sigma_{NN}^{\rm inel}\right]\right]\,.$ (2)
@@
Here, $T_{AB}({\bf b})$ is the nuclear overlap function and $\sigma_{\rm NN}^{\rm inel}$ denotes the nucleon-nucleon inelastic cross-section. For such nucleus-nucleus collisions, the total number of binary collisions is given as:
@@
In heavy-ion physics, theoretical calculations use $\bf b$ as an input to compare theoretical results to the experimental measurements. Both $N_{\rm part}({\bf b})$ and $N_{\rm coll}({\bf b})$ are calculated using Glauber model at a given $\bf b$, which are subsequently related to final state particle multiplicities [90].
In $pp$ collisions, one cannot define impact parameter and centrality as they are point-like collisions of two nucleons. Thus, in experiments, the final state charged particle multiplicity is used to divide the event classes, which we discuss in more detail in the next subsection.
II.3 Charged-particle multiplicity ($N_{\rm ch}$)
@@
One of the most fundamental observables to probe the system size of high-energy collisions is charged-particle multiplicity. In the ALICE experiment at the LHC, the charged-particle multiplicity is estimated at mid- and forward-pseudorapidity regions. The charged-particle multiplicity at the mid-pseudorapidity region (i.e., $|\eta|<0.8$), denoted as $N_{\rm ch}^{\rm mid}$, is measured by the silicon pixel detector (SPD) and the time projection chamber (TPC). Contrary, the measurement of charged-particle multiplicity at the forward pseudorapidity region, denoted as $N_{\rm ch}^{\rm fwd}$, is performed by the V0 (V0A+V0C) detectors having the pseudorapidity coverage of $-3.7<\eta<-1.7$ (V0C) and $5.12.8<\eta<$ 5.1 (V0A). The left and right panels of Fig. 6 show the self-normalised charged-particle multiplicity distributions obtained in mid and forward rapidity, respectively, at $\sqrt{s}=13$ TeV using PYTHIA8. The distributions are further divided into different percentiles and specified as multiplicity classes.
@@
Apart from the SPD, TPC, and V0 detectors discussed above, with the recent upgradation of ALICE in Run 3, the global detector named Fast Interaction Trigger (FIT) is used to estimate the collision centrality and the charged-particle multiplicity. FIT consists of three sub-detectors, namely, FT0, FV0, and FDD. The detailed description of the FIT detector can be found in Ref. [91]. The newly built Muon Forward Tracker (MFT) at forward rapidity ($-3.6<\eta<-2.45$) can also be used to estimate the charged-particle multiplicity in Run 3 of ALICE [92]. Additionally, the CMS experiment at the LHC uses pairs of pixel clusters (known as tracklets) from two different layers (disks) of the silicon pixel detector to measure the charged-particle multiplicity produced in the range $|\eta|<2.6$. These pairs have clusters with relatively small differences in $\eta$ and $\phi$. The correlation between $\eta$ and $\phi$ can be used to select tracklets corresponding to primary charged hadrons. The detailed description of tracklets and vertex reconstruction algorithms can be found in Ref. [93, 94, 95]. The ATLAS experiment at the LHC uses tracking detectors such as the Inner Detector (ID) and the trigger detectors to measure the charged-particle multiplicity. The ID consists of a silicon pixel detector (Pixel), a silicon microstrip detector (SCT), and a transition radiation tracker (TRT) with almost full coverage in $\phi$, and $|\eta|<2.5$. The detailed description of trigger selection and track reconstruction algorithms can be found in Ref. [96, 97, 98]. These three (ALICE, CMS, and ATLAS) are the major experiments at the LHC, and their measurement of final-state charged particle multiplicity as an event classifier is discussed.
II.4 Number of multi-partonic interactions ($N_{\rm mpi}$)
Figure 7: Distribution of number of multi-partonic interactions ($N_{\rm mpi}$) in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8.
@@ II.4 Number of multi-partonic interactions ($N_{\rm mpi}$)
Figure 8: Correlation of mid (left panel) and forward (right panel) rapidity charged-particle multiplicity with number of MPI in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8.
@@
Hadronic collisions at RHIC and LHC produce a large number of particles in the final state. The fundamental description of particle production in hadronic collisions, such as $pp$, can be understood using the QCD-inspired parton model. [99]. In this model, the hadrons are treated as a collection of elementary constituents – quarks and gluons, collectively referred to as partons – bound together by the strong interaction. In minimum-bias hadronic collisions, multiple partonic interactions are inevitable due to the composite structure of hadrons [99, 100], with these MPIs typically being soft in nature. Additionally, in a hard $pp$ collision, which is used for testing the Standard Model (SM) to high accuracy or sometimes to hunt for physics beyond the SM, a hard scattering is accompanied by a number of additional soft interactions, known as underlying event activity (UE). The comprehension and accurate modelling of these additional soft interactions are thus crucial for properly interpreting these hard interactions. Due to the non-perturbative nature, large numbers, and interplay among these interactions, the current approximation of these additional scattering can only be modelled in the context of Monte Carlo (MC). Here, for an event with a fixed final state invariant mass, with the increase in the collision energy of the hadronic species, one can probe to a lower momentum fraction $x$, leading to a higher partonic flux and an increase in the multiple-partonic scattering cross-section [99, 100]. Consequently, events with a lower invariant mass in the final state can have influence from larger multiple hard scatterings [100]. In addition, most inelastic events in hadronic collisions contain several perturbatively calculable interactions when the pQCD theory is extended to the low-$p_{\rm T}$ regime with some distance above $\Lambda_{\rm QCD}$ [99]. Such interactions in inelastic hadronic collisions, even though soft and contributing to MPIs abundantly, are of much significance as they can significantly affect the particle multiplicity in the final state [99]. Thus, with an increase in the soft event activity, one expects the corresponding $N_{\rm mpi}$ to rise. In other words, an event with a large value of $N_{\rm mpi}$ is most likely to be a softer event.
At the LHC, recent measurements of heavy-ion-like signatures for high multiplicity $pp$ collisions can be explained qualitatively with perturbative QCD-based event generators such as PYTHIA8 [19]. These heavy-ion-like features are well reproduced when one probes high MPIs with color reconnection. However, the estimation of $N_{\rm mpi}$ in experiments is difficult. Thus, one looks for event shape observables that can be proportional to $N_{\rm mpi}$. This makes the studies of the correlation of MPI with event-shape observables crucial in MC simulations. Figure 7 shows the distribution of $N_{\rm mpi}$ obtained in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. Different colour representations in the figure denote different percentile cuts of the distribution, specified as different multiplicity classes.
Figure 8 shows the correlation of mid-rapidity charged-particle multiplicity ($N_{\rm ch}^{\rm mid}$) (left panel) and forward rapidity charged-particle multiplicity ($N_{\rm ch}^{\rm fwd}$) (right panel) with the number of $N_{\rm mpi}$ in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. The solid line shows the mean value of $N_{\rm mpi}$, i.e., $\langle N_{\rm mpi}\rangle$, for a given multiplicity value on the $x-$axis. As one notices, both mid- and forward-charged-particle multiplicity retain a fair degree of correlation with $N_{\rm mpi}$. From the figure, it is clear that irrespective of the pseudorapidity region, as the charged-particle multiplicity increases, the corresponding $\langle N_{\rm mpi}\rangle$ also increases, indicating that the events having higher value of $N_{\rm mpi}$ tend to produce a large number of particles at the final state and are softer as compared to the events having lower $N_{\rm mpi}$ value. However, a saturation behavior of the correlation of $N_{\rm mpi}$ with $N_{\rm ch}^{\rm mid}$ is seen for higher values of $N_{\rm ch}^{\rm mid}$. This indicates that beyond $N_{\rm ch}^{\rm mid}\simeq 50$, the increase in multiplicity in the mid-rapidity region probes the events with similar $N_{\rm mpi}$, $i.e,~{}\simeq$ 20. In addition, one notices that using $N_{\rm ch}^{\rm fwd}$ one can reach higher value of $\langle N_{\rm mpi}\rangle$ as compared to the usage of $N_{\rm ch}^{\rm mid}$.
II.5 Transverse sphericity ($S_{\rm{T}}$)
Transverse sphericity, one of the oldest event shape observables, was used at the Stanford Linear Accelerator Center (SLAC) to test the existence of jets in $e^{+}e^{-}$ processes at collision energies up to 7.4 GeV [101]. Transverse sphericity is defined in terms of the eigenvalues of the transverse momentum matrix ($S_{\rm xy}^{\rm Q}$), given as [44]:
$S_{\rm xy}^{\rm Q}=\frac{1}{\sum_{i}p_{{\rm T}_{i}}}\sum_{i}\begin{pmatrix}p_{{\rm x}_{i}}^{2}&p_{{\rm x}_{i}}p_{{\rm y}_{i}}\\ p_{{\rm y}_{i}}p_{{\rm x}_{i}}&p_{{\rm y}_{i}}^{2}\end{pmatrix}\text{[GeV/{\hbox{c}}].}$. (6)
@@ II.5 Transverse sphericity ($S_{\rm{T}}$)
Figure 10: Correlation of $S_{\rm T}$ with number of MPI.
@@
By construction, the value of $S_{\rm T}$ lie between 0 and 1. Here, the extreme limits, i.e., $S_{\rm T}\rightarrow 0$ selects the jetty or pencil-like events, while $S_{\rm T}\rightarrow 1$ indicates that the events are isotropic, which are dominated by the soft production of particles. In this work, $S_{\rm T}$ has been estimated with the charged particles in the mid-pseudorapidity region, i.e., $|\eta|<0.8$ having $p_{\rm T}>0.15$ GeV/$c$. Figure 9 shows the distribution of transverse sphericity measured in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. In addition, different percentile slices are shown in the $S_{\rm T}$ distribution, referred to as sphericity classes, which are shown with different colours. Figure 10 shows the correlation of transverse sphericity with the number of multi-partonic interactions in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8, where the solid line shows the mean value of $N_{\rm mpi}$ for each bin in $S_{\rm T}$. From Fig. 10, it is clear that the transverse sphericity possesses a finite positive correlation with the number of multi-partonic interactions.
II.6 Transverse spherocity ($S_{0}$)
Figure 11: Depiction of isotropic and jetty events based on the transverse spherocity selection [58].

### v2

@@ 2 Event Classifiers
In this section, we define different event shape observables used at the LHC energies. In addition, the correlation of these event shape observables with $N_{\rm mpi}$ is shown and discussed.
2.1 Different types of events: elastic and inelastic collisions
@@
The final state particle production in hadronic collisions, such as proton-on-proton, is a probe for the initial interactions among the colliding hadrons. An interaction can be classified as elastic or inelastic. In elastic collisions, the incoming and outgoing protons are the same, which takes place via exchange of Pomeron [^1] Pomeron is a colour singlet particle that dominates the elastic scatterings at high energies and has the quantum numbers of the vacuum. ALICE:2012fjm . In contrast, in inelastic collisions, the final state particles are not identical to the initial colliding protons. An inelastic event can be diffractive, which is signified by a small energy exchange between the protons. Sometimes, one or both protons interact with the Pomerons and dissociate into multiple final-state particles. The event is single diffractive (SD) if only one proton dissociates while the other one survives as it is ($p+p\rightarrow p+X$). Similarly, when both the protons dissociate to form multiple particles in the final state, it is called a double diffractive (DD) event ($p+p\rightarrow X_{1}+X_{2}$). Sometimes, two Pomerons are exchanged, and both the protons survive along with the production of new particles, i.e., $p+p\rightarrow p+p+X$, it is called a central diffractive (CD) event. Additionally, when more than two Pomerons are exchanged, there is a possibility for both protons to dissociate and lead to the production of new particles. These events are usually accompanied by large particle multiplicity in the final state as compared to other diffractive events. Figure 3 shows the pictorial representation of different types of possible interactions in $pp$ collisions. The inelastic events without SD are called Non-Single Diffractive (NSD). At the LHC, the measurements are usually performed for NSD or inelastic (INEL) events.
2.2 Impact parameter and centrality
@@
Figure 4: A schematic representation of the correlation between the final state charged-particle multiplicity distribution ($N_{\rm ch}$) with the Glauber calculated quantities ($b$, and $N_{\rm part}$). The collision centrality classification in experiments is shown explicitly. Figure is taken from Ref. Miller:2007ri
@@
The impact parameter ($b$) of a collision is one of the key parameters that substantially influence the particle production in the final state. However, its measurement is not trivial in experiments as the length scale of the impact parameter ranges in the level of a few fermis. Thus, experimental data are typically categorised by the centrality, defined as percentile of events obtained from the number of produced particles or by the number of participant nucleons, $N_{\rm part}$ that is registered in detectors through the estimation of number of spectator nucleons via $N_{\rm part}=2A-N_{\rm spec}$, where $A$ is the mass number of the colliding nuclei (for symmetric collisions) and $N_{\rm spec}$ is the number of spectator nucleons. For example, in ALICE experiment, $N_{\rm spec}$ is given by $E_{\rm ZDC}/E_{\rm A}$, where, $E_{\rm ZDC}$ is the energy deposited in zero-degree calorimeters (ZDC) and $E_{\rm A}$ is beam energy per nucleon. Centrality allows for dividing events into different classes. However, phenomenologically, one needs to assign an impact parameter to a given centrality as depicted in Fig. 4. Thus, theoretical techniques, using the Glauber formalism Miller:2007ri , Loizides:2016 , Glauber:1970 , Wong:1994book have been developed to allow the estimation of impact parameter along with the number of participant nucleons from experimental data, which considers multiple scattering of nucleons in nuclear targets. Here, we briefly explain how the total inelastic cross section, the number of binary collisions, and the number of participants depend on the impact parameter.
@@
Figure 5: Schematic representation of the Optical Glauber Model geometry, with transverse (a) and longitudinal (b) views. Figure is taken from Ref. Miller:2007ri
@@
For a collision of two heavy-nuclei, $A$ and $B$ at relativistic speeds with impact parameter $\bf b$} as shown in Fig. 5, the inelastic cross section can be defined as
$\sigma^{\rm inel}_{AB}({\bf b})$ $=$ $\int 2\pi b\;d{\bf b}\left[1-\left(1-T_{AB}({\bf b})\sigma_{NN}^{\rm inel}\right)^{AB}\right]$ (1)
$\simeq$ $\int 2\pi b\;\;d{\bf b}\left[1-\exp\left[-ABT_{AB}({\bf b})\sigma_{NN}^{\rm inel}\right]\right]\,.$ (2)
@@
Here, $T_{AB}({\bf b})$ is the nuclear overlap function and $\sigma_{\rm NN}^{\rm inel}$ denotes the nucleon-nucleon inelastic cross section. For such nucleus-nucleus collisions, the total number of binary collisions is given as:
@@
In heavy-ion physics, theoretical calculations use $\bf b$ as an input to compare theoretical results to the experimental measurements. Both $N_{\rm part}({\bf b})$ and $N_{\rm coll}({\bf b})$ are calculated using Glauber model at a given $\bf b$, which are subsequently related to final state particle multiplicities Kolb:2001 .
In $pp$ collisions, one cannot define impact parameter and centrality as they are point-like collisions of two nucleons. In the Glauber model framework, which is used to describe heavy-ion collisions, treating the proton as point-like is an approximation. Thus, in experiments, the final state charged particle multiplicity is used to divide the event classes, which we discuss in more detail in the next subsection.
2.3 Charged-particle multiplicity ($N_{\rm ch}$)
@@
One of the most fundamental observables to probe the system size of high-energy collisions is charged-particle multiplicity. In the ALICE experiment at the LHC, the charged-particle multiplicity is estimated at mid- and forward-pseudorapidity regions. The charged-particle multiplicity at the mid-pseudorapidity region (i.e., $|\eta|<0.8$), denoted as $N_{\rm ch}^{\rm mid}$, is measured by the silicon pixel detector (SPD) and the time projection chamber (TPC). Contrary, the measurement of charged-particle multiplicity at the forward pseudorapidity region, denoted as $N_{\rm ch}^{\rm fwd}$, is performed by the V0 (V0A+V0C) detectors having the pseudorapidity coverage of $3.7<\eta<-1.7$ (V0C) and $2.8<\eta<5.1$ (V0A). The left and right panels of Fig. 6 show the self-normalised charged-particle multiplicity distributions obtained in mid and forward rapidity, respectively, at $\sqrt{s}=13$ TeV using PYTHIA8. The distributions are further divided into different percentiles and specified as multiplicity classes.
@@
Apart from the SPD, TPC, and V0 detectors discussed above, with the recent upgradation of ALICE in Run 3, the global detector named Fast Interaction Trigger (FIT) is used to estimate the collision centrality and the charged-particle multiplicity. FIT consists of three sub-detectors, namely, FT0, FV0, and FDD. The detailed description of the FIT detector can be found in Ref. Bysiak:FIT . The newly built Muon Forward Tracker (MFT) at forward rapidity ($3.6<\eta<-2.5$) can also be used to estimate the charged-particle multiplicity in Run 3 of ALICE ALICE:MFTTDR . Additionally, the CMS experiment at the LHC uses pairs of pixel clusters (known as tracklets) from two different layers (disks) of the silicon pixel detector to measure the charged-particle multiplicity produced in the range $|\eta|<2.6$. These pairs have clusters with relatively small differences in $\eta$ and $\phi$. The correlation between $\eta$ and $\phi$ can be used to select tracklets corresponding to primary charged hadrons. The detailed description of tracklets and vertex reconstruction algorithms can be found in Ref. CMS:2024ykx , CMS:2011aqh , CMS:2010qvf . The ATLAS experiment at the LHC uses tracking detectors such as the Inner Detector (ID) and the trigger detectors to measure the charged-particle multiplicity. The ID consists of a silicon pixel detector (Pixel), a silicon microstrip detector (SCT), and a transition radiation tracker (TRT) with almost full coverage in $\phi$, and $|\eta|<2.5$. The detailed description of trigger selection and track reconstruction algorithms can be found in Ref. ATLAS:2010jvh , ATLAS:2010zmc , ATLAS:2016zkp . These three (ALICE, CMS, and ATLAS) are the major experiments at the LHC, and their measurement of final-state charged particle multiplicity as an event classifier is discussed.
2.4 Number of multi-partonic interactions ($N_{\rm mpi}$)
Figure 7: Distribution of $N_{\rm mpi}$ in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8.
Hadronic collisions at RHIC and LHC produce a large number of particles in the final state. The fundamental description of particle production in hadronic collisions, such as $pp$, can be understood using the QCD-inspired parton model. Sjostrand:2004pf . In this model, the hadrons are treated as a collection of elementary constituents – quarks and gluons, collectively referred to as partons – bound together by the strong interaction. In minimum-bias hadronic collisions, multiple partonic interactions are inevitable due to the composite structure of hadrons Sjostrand:2004pf , Bartalini:2011jp , with these MPIs typically being soft in nature. Additionally, in a hard $pp$ collision, which is used for testing the Standard Model (SM) to high accuracy or sometimes to hunt for physics beyond the SM, a hard scattering is accompanied by a number of additional soft interactions, known as underlying event activity (UE). The comprehension and accurate modelling of these additional soft interactions are thus crucial for properly interpreting these hard interactions. Due to the non-perturbative nature, large numbers, and interplay among these interactions, the current approximation of these additional scattering can only be modelled in the context of Monte Carlo (MC) simulation. Here, for an event with a fixed final state invariant mass, with the increase in the collision energy of the hadronic species, one can probe to a lower momentum fraction $x$, leading to a higher partonic flux and an increase in the multiple-partonic scattering cross section Sjostrand:2004pf , Bartalini:2011jp . Consequently, events with a lower invariant mass in the final state can have influence from larger multiple hard scatterings Bartalini:2011jp . In addition, most inelastic events in hadronic collisions contain several perturbatively calculable interactions when the pQCD theory is extended to the low-$p_{\rm T}$ regime with some distance above $\Lambda_{\rm QCD}$ Sjostrand:2004pf . Such interactions in inelastic hadronic collisions, even though soft and contributing to MPIs abundantly, are of much significance as they can significantly affect the particle multiplicity in the final state Sjostrand:2004pf . Thus, with an increase in the soft event activity, one expects the corresponding $N_{\rm mpi}$ to rise. In other words, an event with a large value of $N_{\rm mpi}$ is most likely to be a softer event.
The authors in Ref. Christiansen:2024bhe provide an abstract discussion of MPI that can be applicable to most of the models. As MPIs are modeled as independent $2\rightarrow 2$ partonic interactions, the corresponding pQCD cross section diverges and finally exceeds the total inelastic cross section for the limit $p_{\rm T}\rightarrow 0$. Since the ratio of pQCD cross section to the total cross section is interpreted as the average number of MPIs per event, the divergence is not inherently problematic. Nevertheless, to avoid nonphysical values of cross sections at $p_{\rm T}\rightarrow 0$, regularization is necessary. This regularization can be associated with color-screening effects, which limit the spatial resolution of individual partons as implemented in PYTHIA, or with initial-state saturation mechanisms as in EPOS.
In PYTHIA, a transverse momentum cut off scale, $p_{\rm T0}$ is introduced to regularize $2\rightarrow 2$ partonic subprocesses, as follows:
$\frac{{\rm d}\hat{\sigma}}{{\rm d}p_{\rm T}^{2}}\propto\frac{\alpha_{S}^{2}(p_{\rm T}^{2})}{p_{\rm T}^{4}}\rightarrow\frac{\alpha_{S}^{2}(p_{\rm T0}^{2}+p_{\rm T}^{2})}{(p_{\rm T0}^{2}+p_{\rm T}^{2})^{2}}.$ (6)
Here, $p_{\rm T0}$ depends upon the center of mass energy of the collision, given as:
$p_{\rm T_{0}}(s)=p_{\rm T}^{\rm ref}\Bigg(\frac{\sqrt{s}}{\sqrt{s_{0}}}\Bigg)^{\epsilon}.$ (7)
Here, $p_{\rm T}^{\rm ref}$ and $\epsilon$ are free parameters that can be tuned to the experiments. $p_{\rm T0}$ increases with an increase in collision energy, which is due to the increase in the low-$x$ partons in the wavefunction of colliding protons. In PYTHIA 8, $N_{\rm{mpi}}$ additionally depends on the impact parameter, which is determined by the spatial overlap of the transverse profiles of the colliding protons. This impact-parameter dependence introduces a geometrical component to the MPI framework, which makes the description of MPI more realistic.
At the LHC, recent measurements of heavy-ion-like signatures for high multiplicity $pp$ collisions can be explained qualitatively with perturbative QCD-based event generators such as PYTHIA8 ALICE:2019avo . These heavy-ion-like features are well reproduced when one probes high MPIs with color reconnection. However, the estimation of $N_{\rm mpi}$ in experiments is difficult. Thus, one looks for event shape observables that can be proportional to $N_{\rm mpi}$. This makes the studies of the correlation of MPI with event-shape observables crucial in MC simulations. Figure 7 shows the distribution of $N_{\rm mpi}$ obtained in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. Different colour representations in the figure denote different percentile cuts of the distribution, specified as different $N_{\rm mpi}$ classes.
@@ 2.4 Number of multi-partonic interactions ($N_{\rm mpi}$)
Figure 8: Correlation of mid (left panel) and forward (right panel) rapidity charged-particle multiplicity with number of MPI in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. The solid line represents the average value of $y$-axis, i.e., $\langle N_{\rm mpi}\rangle$, for each bin in the $x$-axis, i.e., $N_{\rm ch}^{\rm mid}$ (left) and $N_{\rm ch}^{\rm fwd}$ (right).
[…]
2.5 Transverse sphericity ($S_{\rm{T}}$)
@@
Transverse sphericity, one of the oldest event shape observables, was used at the Stanford Linear Accelerator Center (SLAC) to test the existence of jets in $e^{+}e^{-}$ processes at collision energies up to 7.4 GeV Hanson:1975fe . Transverse sphericity is defined in terms of the eigenvalues of the transverse momentum matrix ($S_{\rm xy}^{\rm Q}$), given as ALICE:2012cor :
$S_{\rm xy}^{\rm Q}=\frac{1}{\sum_{i}p_{{\rm T}_{i}}}\sum_{i}\begin{pmatrix}p_{{\rm x}_{i}}^{2}&p_{{\rm x}_{i}}p_{{\rm y}_{i}}\\ p_{{\rm y}_{i}}p_{{\rm x}_{i}}&p_{{\rm y}_{i}}^{2}\end{pmatrix}\text{[GeV/$c$].}$ (8)
@@
Figure 10: Correlation of $S_{\rm T}$ with number of MPI. The black solid line represents $\langle N_{\rm mpi}\rangle$ as a function of $S_{\rm T}$.
@@
By construction, the value of $S_{\rm T}$ lies between 0 and 1. Here, the extreme limits, i.e., $S_{\rm T}\rightarrow 0$ selects the jetty or pencil-like events, while $S_{\rm T}\rightarrow 1$ indicates that the events are isotropic, which are dominated by the soft production of particles. In this work, $S_{\rm T}$ has been estimated with the charged particles in the mid-pseudorapidity region, i.e., $|\eta|<0.8$ having $p_{\rm T}>0.15$ GeV/$c$. Figure 9 shows the distribution of transverse sphericity measured in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8. In addition, different percentile slices are shown in the $S_{\rm T}$ distribution, referred to as sphericity classes, which are shown with different colours. Figure 10 shows the correlation of transverse sphericity with the number of multi-partonic interactions in $pp$ collisions at $\sqrt{s}=13$ TeV using PYTHIA8, where the solid line shows the mean value of $N_{\rm mpi}$ for each bin in $S_{\rm T}$. From Fig. 10, it is clear that the transverse sphericity possesses a finite positive correlation with the number of multi-partonic interactions.
2.6 Transverse spherocity ($S_{0}$)
Figure 11: Depiction of isotropic and jetty events based on the transverse spherocity selection Khuntia:2018qox .